OpenStudy (idealist10):

Solve (x+y)^2 dx+(x+y)^2 dy=0.

OpenStudy (idealist10):

integral of (x+y)^2 dx=(x+y)^3/3+g(y)=C (x+y)^2+g'(y)=(x+y)^2 g'(y)=0 g(y)=0 (x+y)^3/3=C But the answer in the book is x+y=C. So which answer is right?

OpenStudy (idealist10):

@amistre64

OpenStudy (amistre64):

g(y) = C since g'(y) = 0 so the solution is: (x+y)^3/3 + C is the family of solutions

OpenStudy (idealist10):

Thank you!

OpenStudy (amistre64):

youre welcome